OpenStudy (anonymous):

A rope is square root 250 units long, the rope is cut into two pieces, so that the lengths of the pieces are in the ratio 3:2. what is the length of the longer piece expresses simplest radical form

OpenStudy (anonymous):

A. \[3 \sqrt{10}\] units B. \[10\sqrt{5}\] C. \[10\sqrt{2}\] D. \[2\sqrt{10}\]

OpenStudy (anonymous):

@e.mccormick @CGGURUMANJUNATH @Luigi0210 @Kfins99 @Gin_Ichimaru @grant330sims @hartnn @ganeshie8 help please if you know

OpenStudy (cggurumanjunath):